solve the system of equations
3x+y=3 x+2y=-4

Answer by checkley71(8403)

3x+y=3 or y=-3x+3
x+2(-3x+3)=-4
x-6x+6=-4
-5x=-4-6
-5x=-10
x=-10/-5
x=2 answer.
y=-3*2+3
y=-6+3
y=-3 answer.
proof
2+2*-3=-4
2-6=-4
-4=-4